Home‎ > ‎Reports‎ > ‎Crew 109‎ > ‎

day14


Crew #109 Log Book for December 30, 2011

Photos of the day





Commander's Check-In Report

Date and time: 30 december 2011, 18:30

Written by Nicky De Munster, Crew 109 Commander

Crew Physical Status: Good

Time Departed/Returned from EVA:

PED EVA 22 (Leandro, David, Karon): 9:51 - 12:30
EVA 23 (Nicky, Aster, Victoria): 14:05 - 17:00

Brief Narative of Field Mission Results:

EVA 22: explored Skyline Rim
EVA 23: explored Candor Chasma

EVA Data/Interpretations:

no samples were collected

Engineering/Hab Maintenance:

all systems OK thanks to Leandro!

Report Transmission Schedule:

CDR check-in, CDR, Science (Biology, Geology), EVA, Journalists

Plans for Tomorrow:

Clean the Hab
Crew change-over

Inventory:

Support Requested:

Miscellaneous:

Commander's Report

Date and time: 30 december 2011, 19:00

Written by Nicky De Munster,Crew 109, Commander:

Today we did two last EVA's. Just some exploring by feet to Skyline Rim and a amazing hike to Candor chasma.
While Leandro, David and Karon were EVA this morning, Aster and myself recorded the second part of the science experiments for Eductional Innovations in the lab.
Candor Chasma was a real discovery. Perhaps one of the most beautiful places here.
This afternoon team one cleaned up the lab. Kitchen will be cleaned up tomorrow morning.
Tonight we stopped simulation. Mission accomplished.
Thanks to my amazing crew we did a successful mission and performed 90% of our objectives, which was a great success here at MDRS


Science Report (Geology)

Date and Time: 30 December 2011, 15:59

Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

All soil sites have been hit for a total of 33 cone penetrometer blowcount tests and 11 8-inch core samples. This data will be integrated to produce a map detailing the most effective rover routes. The primary function of these maps is to show trafficability. For example the maps will show: 1) the extent where a rover can travel without a significant risk of flipping over due to an excessively high gradient of the slope it is climbing or descending; 2) the places of loose soil in which a rover can get stuck; 3) places that the soil strength is sufficient to prevent a localized rockslide which could flip or destroy the rover; 4) water content of the soils; 5) Hydrous mineral deposit locations based on a reflectance spectra with absorption minima at 2.2 μm.


Science Report (Biology)

Date and Time: 30 December 2011, 18:00

Written by Karon Wynne, Health & and Safety Office and Crew 109 Biologist

On our last day at MDRS I prepared the Biology lab for the next crew. I carefully disposed of all of our samples from the mission and sterilized the oven. I also wiped down the counter tops and organized all of the cabinets and drawers. The lab now looks great and is ready for crew 110. Crew 109 was able to successfully culture bacteria from samples collected around MDRS and we hope that the next crew will have as much success and perhaps be able to elaborate on our results in the coming weeks.

EVA 22 Report

Date and Time: 30 December 2011, 16:15

Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

Time: 9:51 – 12:30

Crew Members: David, Karon, Leandro

Site Location: Base of Skyline Rim

Transit Mode: On foot

Objectives: Explore Southwestern extent of the base of Skyline Rim

Results: Successful traverse to Base of Skyline Rim

Route spanned 8.6 km.

Lessons learned: None

EVA 23 Report

Date and Time: 30 December 2011 17:00

Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

Time: 15:00 – 17:00

Crew Members: Nicky, Aster, Victoria

Site Location: Candor Chasma

Transit Mode: ATV, on foot

Objectives: Explore Candor Chasma

Results: Successful exploration of Candor Chasma

Route spanned 8.8 km.

Lessons learned: None

Engineering Report

Date and time: 30 December 2011, 18:34 MST

Written by Leandro G. Barajas, Crew 109, Chief Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): 0.75
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75
Notes/Comments: 

Oil:
ATV Oil Quantity (# of Quarts in storage): 0
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 12
Trailer Potable Water Tank Level (inches from bottom): 15
Hab Potable Water Tank Level (inches from bottom): 12
Potable Water Meter Reading: 48366.3
In to GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A
Greenhab Notes/Comments: N/A

Observatory:
Telescope Used? (Yes/No): NO
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: 125475.5
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 0.3 Gallons
Tire Status: Good - 4 psi
Today's Run Time: 00:30
Season's Runtime: 40:40
Comments:

ATV 2 (Spirit):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 0.3 Gallon
Tire Status: Good - 4 psi
Today's Run Time: 00:30
Season's Run Time: 34:14
Comments:

ATV 3 (Viking I):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 0.3 Gallons
Tire Status: Good - 4 psi
Today's Run Time: 00:30
Season's Run Time 37:25
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Fahrenheit): 60 
Upstairs Furnace has not been on since the downstairs furnace was enabled.
Actual temperature upstairs (Fahrenheit): 75

Thermostat setting downstairs (Fahrenheit): 55 
Thermostat is not operational. Currently using a toggle switch to turn on furnace a few hours a day (Not used today).
Actual temperature downstairs (Fahrenheit): 65

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering Notes/Comments:
Heat Tapes Operating 24/7 (Both Lights On).

The following are engineering items that should be passed to the next MDRS crew for continuity and should be addressed by the responsible party in MDRS if necessary:
- There is a lack of spare parts to repair the suits; specifically connectors and fittings and power switches.
There is some hose but we should have more.
- There is no ATV oil in storage. Next crew most like will need it.
- There's no spare oil for the Hab Car
- ATV 3 need to get a replacement Taillight Lens
- ATV 2 seems to have the idle nozzle partially blocked.
It requires many tries to cold start the ATV and it turns off frequently when in idle.
- There is no operational downstairs thermostat. Currently a toggle switch is used to control the furnace. Either the thermostat or the wiring seems to be defective.
- The observatory computer may turn off randomly. Someone will need to go into the observatory and restart it.
- There are only 3 radios fully operational. 1 radio, only works in receive mode, it does not transmit. The other 2 radios do not work with the original rechargeable batteries, however, they seem to work with non-rechargable batteries, or with standard AA rechargable batteries charged on a external battery charger.
- There are 4 radio headsets that work with the radios, but only in VOX mode. If the PTT is pressed, then the radio has to be power-cycled in order to get the headset to work again.
- Radios are not really usable while riding the ATVs, however, if they are set at full volume, clicking the PTT is enough to call the attention of the other EVA crew members.
- Two fluorescent lights on the first floor need to be replaced.

Journalist Report

Date and Time: 30 December 2011

Written by: Aster Stein, Crew 109, Journalist

Here it is, my final report as Crew 109 journalist. This was our last full day at the Hab, the two past weeks have gone by so fast. Everyone here is happy to go back home, but there is of course a tiny part of sadness that we have to leave our hab behind, in the hands of complete strangers! We were only now beginning to understand how to run a remote Mars base.

Then again, we were those strangers two weeks ago. And we took pretty good care of things here. Crew 110 will surely do the same thing. There will be a lot to take in on their first days though. Where is everything? How does everything work? What are those things scratching at your door at night? What do you mean, don’t add too much love to the breads? Before leaving the airlock one last time, we’ll bestow as much Hab lore upon them, and then they’re on their own on their own adventure.

Crew 109 was a unique group of people. Our paths crossed, out of all places, in the middle of nowhere in the desert in Utah. Each one of us has their own quirks and personality, but it all worked out. There were no fights, no arguments, if my information is accurate, not even disagreements. We have lived and worked together for two weeks and got to know each other a bit. It just all happened naturally. We had a lot of fun learning and exploring Mars together. Problems? We took ‘em and threw ‘em to the ground!

I would like to thank Karon, Nicky, Leandro and Victoria for being such awesome crew mates. I would also like to thank David for the same, but he gets a tiny extra thanks for writing our daily haiku’s. It’s been a privilege making all those happy memories together.

But now we have to part ways again, and wake up from our two week dream at the the hab. Back to real life. Odds are that we will not meet each other again. Then again, who knows… Until we do, we’ll always have Mars.

Rocks, robots, balloons

Completely schooled this planet

Acquired friendships too


This is crew 109, signing off.


Chef's Report

Date and time: 30 December 2011 18:30

Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

1a) Today was
- cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) or activities were prepared to mark the celebration?
No 

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Gumbo + ground beef and chocolate mousse

3) List main meal foods not finished by the crew (leftovers stored for later use or discarded)

None

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Chocolate mousse:
6 egg whites and 6 egg yolks (powder equivilant)
150 grams powdered sugar
200 grams pure chocolate

- any cooking questions you need help with?
No

- any unusual events? (cleaned the oven, spilled all the vanilla, took inventory)
No

- any other thoughts or suggestions?
No